# Applying the Spectral Method

### v1

@@
Abstract. This paper proposes a new technique for computer modeling linear filters based on the spectral form of mathematical description of linear systems. It assumes that input and output signals of the filter are represented as orthogonal expansions, while filters themselves are described by two-dimensional non-stationary transfer functions. This technique allows one to model the output signal in continuous time, and it is successfully tested on the Butterworth, Linkwitz–Riley, and Chebyshev filters with different orders.
@@ 1. Introduction
The paper concerns the problem of signal processing by linear filters. A linear filter is a linear dynamic system that passes components of the input signal with frequencies from a certain range (passband) and suppresses components of the input signal with frequencies that do not belong to such a range (stopband) [1].
@@
This field began to actively develop about a hundred years ago. The theory of linear filters is inextricably linked with the research of Ronald Martin Foster [3], Wilhelm Cauer [4], Stephen Butterworth [5], Hendrik Wade Bode [6], Sidney Darlington [7], and others. The early stages of this theory are thoroughly described in reviews [8, 9].
@@
As a rule, the transition to discrete time is used for computer modeling linear filters [15, 26, 27, 28]. However, it is possible to propose a modeling technique that does not require time discretization (continuous 3time modeling is more natural for analog filters). This technique is based on the spectral form of mathematical description of linear systems [29, 30, 31], which is oriented towards computer modeling from the moment of its appearance. The purpose of the paper is to describe and test the spectral method for modeling linear filters such as the Butterworth, Linkwitz–Riley, and Chebyshev filters. In this paper, the Linkwitz–Riley filters are considered to show how the spectral form is applied to describe a cascade of linear filters.
@@ 2. Description of Linear Filters and Problem Statement
In this section and below, we use standard notations and terms from complex analysis [32]. In particular, $\mathds{C}$ represents the set of complex numbers (the complex plane); $\mathrm{i}$ is the imaginary unit; $\mathop{\mathrm{Re}}z$ and $\mathop{\mathrm{Im}}z$ denote the real and imaginary parts of a number $z\in\mathds{C}$ which give the Cartesian coordinates of a point in the complex plane; $|z|$ and $\arg z$ are its absolute value and argument which determine the polar coordinates of a point on the complex plane; $\bar{z}$ is the complex conjugate of $z$. The sets $\{z\colon\mathop{\mathrm{Re}}z<0\}$ and $\{z\colon\mathop{\mathrm{Im}}z>0\}$ are the left and upper half-planes, respectively. In addition, some definitions and notations from the theory of function spaces are used, namely spaces of square-integrable functions $L_{2}(\mathds{T})$} and $L_{2}(\mathds{T}^{2})$}, as well as their bases [33].
@@ 1. Description by linear differential equations:
where $g(t)$ and $x(t)$ are input and output signals, respectively, $a_{0},a_{1},\dots,a_{n}$ and $b_{0},b_{1},\dots,b_{m}$ are given values, $n$ represents the filter order. The initial conditions are assumed to be zero.
@@ 4. Description by spectral transforms:
The matrix $W$ can be expressed by coefficients $a_{0},a_{1},\dots,a_{n}$ and $b_{0},b_{1},\dots,b_{m}$:
@@ 3. The Spectral Form of Mathematical Description
1. The term “non-stationary” is used when describing signals and linear systems in the spectral form. In [29], as well as in earlier and later publications on the spectral theory of control systems, the non-stationarity is understood in the following sense: boundaries of the time segment, on which a linear system should be analyzed change, over time (both boundaries may change, or only one of them, usually the right boundary), i.e., the time segment is non-stationary. In this paper, the time segment is assumed to be fixed, or stationary.
2. Modern publications on the spectral form of mathematical description use different terminology. For example, in [34, 35], the presentation is in terms of spectral characteristics of functions and linear operators. However, the results of this paper relate to signal processing, and the classical terminology from [29] is more preferable in this context.
3. Despite the fact that when describing linear systems by integral and spectral transforms, signals and a linear system itself are represented differently (as functions of a complex variable in the first case and as infinite column matrices and infinite matrices in the second case), these forms have similar properties. Both forms facilitate the transition from differential and integral relations to algebraic relations between input and output signals. Furthermore, for both forms, characteristics of a linear system are easily expressed through coefficients of the left-hand and right-hand sides of the equation (1).
4. The formulae (4) and (7) have the same structure that allows one to easily express the two-dimensional non-stationary transfer function by the transfer function and vice versa. In addition, the representation of the transfer function
$H(s)=\frac{b_{m}(s-\nu_{1})(s-\nu_{2})\cdots(s-\nu_{m})}{a_{n}(s-s_{1})(s-s_{2})\cdots(s-s_{n})},$ (8)
@@
$W=\frac{b_{m}}{a_{n}}\,(P-s_{1}E)^{-1}(P-s_{2}E)^{-1}\cdots(P-s_{n}E)^{-1}(P-\nu_{1}E)(P-\nu_{2}E)\cdots(P-\nu_{m}E).$ (9)
5. The domain for the spectral transform of functions of one variable is the space $L_{2}(\mathds{T})$}. However, the spectral transform can also be applied to generalized functions [29], as well as random processes, including generalized random processes of the white noise type [31]. For the listed variants, the result of the spectral transform is the non-stationary spectral characteristic. The domain for the spectral transform of functions of two variables is also not limited to the space $L_{2}(\mathds{T}^{2})$}, and generalized functions are included in it. Four simplest linear systems are considered below, and only one of them corresponds to an impulse transfer function from $L_{2}(\mathds{T}^{2})$}; otherwise, impulse transfer functions are generalized functions, and two-dimensional non-stationary transfer functions are defined for all linear systems.
6. When analyzing and synthesizing linear filters, frequency characteristics determined by the Fourier transform are used first. The spectral transform is associated with generalized Fourier series, and arbitrary orthonormal bases can be used to represent signals. Below, trigonometric functions are used as a basis, but other bases can be used with the same success, e.g., orthogonal polynomials, piecewise constant functions, etc.
@@
Elements of the matrix $A$ (for the proportional block with gain $a(t)$) are calculated using the formula
$A_{ij}=\int_{\mathds{T}}a(t)q(i,t)q(j,t)dt,\ \ \ i,j=0,1,2,\dots,$
elements of matrices $P^{-1}$ and $P$ (for integral and derivative blocks) satisfy the following relations:
@@
where $\tau<0$ corresponds to a time delay, and $\tau>0$ corresponds to a time advance. As a rule, in the latter formula, basis functions $q(i,t)$ are extended by zero: $q(i,t+\tau)=0$ for $t+\tau\notin\mathds{T}$, $i=0,1,2,\dots$ However, we redefine them differently (this is described in more detail below). In the problem under consideration, this improves the quality of modeling linear filters in examples.
@@
forming the orthonormal basis of $L_{2}(\mathds{T})$}, are given below. In these formulae, the indices take the following values: $i,j=1,2,\dots$ and $i\neq j$.
@@
Formally, the domain of basis functions is the segment $\mathds{T}$, but in this case, basis functions can be defined beyond $\mathds{T}$ in a natural way, i.e., values of basis functions beyond $\mathds{T}$ are determined by the same formula as on $\mathds{T}$. Such a definition is used for calculating elements of the matrix $S$, and another method is discussed in Appendix.
@@
Elements of the column matrix $F$ corresponding to the signal $f(t)=\sin\omega t$ are
$F_{0}=\frac{1-\cos T\omega}{\sqrt{T}\omega},\ \ \ F_{i}=\frac{T\sqrt{2T}\omega[(-1)^{i}\cos T\omega-1]}{i^{2}\pi^{2}-T^{2}\omega^{2}}\ \ \ \text{for}\ \ \ i\pi\neq T\omega,\ \ \ F_{i}=0\ \ \ \text{for}\ \ \ i\pi=T\omega,$
and elements of the column matrix $F$ corresponding to the signal $f(t)=\cos\omega t$ are
@@ 4.1. The Butterworth Filters
The domain coloring technique [32] can be used to visualize the characteristic polynomial of the Butterworth filter (the transfer function denominator), see Figure 1. Here, a version of this technique based on the HLS color model is applied, i.e.,
@@
Figure 1 shows a fragment of the complex plane: $\{z\colon\mathop{\mathrm{Re}}z,\mathop{\mathrm{Im}}z\in[-2,2]\}$, the filter order is $n=3$. Zeros of the characteristic polynomial (transfer function poles) correspond to three points on the unit circle: 1-1, $(-1\pm\mathrm{i}\sqrt{3})/2$.
@@ 4.3. The Chebyshev Type I Filters
A visualization for the characteristic polynomial of the Chebyshev Type I filter (the transfer function denominator) is carried out in the same way as for the Butterworth filter; the filter order is $n=5$, see Figure 3. Zeros of the characteristic polynomial (transfer function poles) correspond to five points; they lie on the ellipse that can be obtained by scaling the unit circle along the axes.
@@ 4.4. The Chebyshev Type II Filters
The transfer function of the $n$th-order Chebyshev Type II filter is specified using previously defined parameters (in this case, the parameter $\varepsilon>0$ that determines the constant $\gamma$ is the stopband ripple factor), namely
$H^{\scriptscriptstyle\mathrm{CII}}(s)=\frac{\gamma}{(s-s_{1}^{\scriptscriptstyle\mathrm{CII}})\dots(s-s_{n}^{\scriptscriptstyle\mathrm{CII}})}=\gamma\biggl{[}\,\prod\limits_{k=1}^{n}(s-s_{k}^{\scriptscriptstyle\mathrm{CII}})\biggr{]}^{-1},$ (23)
@@
$s_{k}^{\scriptscriptstyle\mathrm{CII}}=\frac{1}{s_{k}^{\scriptscriptstyle\mathrm{CI}}}=\frac{\bar{s}_{k}^{\scriptscriptstyle\mathrm{CI}}}{|s_{k}^{\scriptscriptstyle\mathrm{CI}}|^{2}},\ \ \ k=1,\dots,n.$ (24)
The transfer function poles can be defined as roots of the equation $\varepsilon^{2}T_{n}^{2}(1/z)=-1$, belonging to the upper half-plane, with the coefficient $\mathrm{i}$. Using relations (11), (19), and (24), we have
@@
$W^{\scriptscriptstyle\mathrm{CII}}(P)=\gamma\bigl{(}(P-s_{1}^{\scriptscriptstyle\mathrm{CII}}E)\dots(P-s_{n}^{\scriptscriptstyle\mathrm{CII}}E)\bigr{)}^{-1}=\gamma\biggl{[}\,\prod\limits_{k=1}^{n}(P-s_{k}^{\scriptscriptstyle\mathrm{CII}}E)\biggr{]}^{-1}.$ (26)
@@
and $(s-s_{k}^{\scriptscriptstyle\mathrm{CII}})(s-s_{n-k+1}^{\scriptscriptstyle\mathrm{CII}})=s^{2}+2\alpha A_{k}s/\rho_{k}^{2}+1/\rho_{k}^{2}$. This yields the expression equivalent to the formula (23):
$H^{\scriptscriptstyle\mathrm{CII}}(s)=\left\{\begin{array}[]{ll}\displaystyle\gamma\biggl{[}\,\prod\limits_{k=1}^{n/2}\biggl{(}s^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,s+\frac{1}{\rho_{k}^{2}}\biggr{)}\biggr{]}^{-1}&\text{for even $n$}\\ \displaystyle\gamma\biggl{[}\biggl{(}s+\frac{1}{\alpha}\biggr{)}\prod\limits_{k=1}^{(n-1)/2}\biggl{(}s^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,s+\frac{1}{\rho_{k}^{2}}\biggr{)}\biggr{]}^{-1}&\text{for odd $n$},\end{array}\right$. (27)
@@
$W^{\scriptscriptstyle\mathrm{CII}}(P)=\left\{\begin{array}[]{ll}\displaystyle\gamma\biggl{[}\,\prod\limits_{k=1}^{n/2}\biggl{(}P^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,P+\frac{1}{\rho_{k}^{2}}\,E\biggr{)}\biggr{]}^{-1}&\text{for even $n$}\\ \displaystyle\gamma\biggl{[}\biggl{(}P+\frac{1}{\alpha}\,E\biggr{)}\prod\limits_{k=1}^{(n-1)/2}\biggl{(}P^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,P+\frac{1}{\rho_{k}^{2}}\,E\biggr{)}\biggr{]}^{-1}&\text{for odd $n$}.\end{array}\right$. (28)

### v2

@@
Abstract. This paper proposes a new technique for computer modeling linear filters based on the spectral form of mathematical description of linear systems. It assumes the representation of input and output signals of the filter as orthogonal expansions, while filters themselves are described by two-dimensional non-stationary transfer functions. This technique allows one to model the output signal in continuous time, and it is successfully tested on the Butterworth, Linkwitz–Riley, and Chebyshev filters with different orders.
@@
00footnotetext: Citation: Rybakov, K.A.; Shermatov, E.D. Applying the spectral method for modeling linear filters: Butterworth, Linkwitz–Riley, and Chebyshev filters. Franklin Open 2026, 14, 100508. https://doi.org/10.1016/j.fraope.2026.100508
@@ 1. Introduction
The paper concerns the problem of signal processing by linear filters. A linear filter is a linear dynamic system that passes components of the input signal with frequencies from a certain range (passband) and suppresses components of the input signal with frequencies that do not belong to such a range (stopband) [1, 2].
@@
This field began to actively develop about a hundred years ago. The theory of linear filters is inextricably related to the research of Ronald Martin Foster [4], Wilhelm Cauer [5], Stephen Butterworth [6], Hendrik Wade Bode [7], Sidney Darlington [8], and others. The early stages of this theory are thoroughly described in reviews [9, 10].
@@
As a rule, the transition to discrete time is used for computer modeling linear filters [27, 28, 29, 30]. However, it is possible to propose a modeling technique that does not require time discretization (continuous time modeling is more natural for analog filters). This technique is based on the spectral form of mathematical description of linear systems [31, 32, 33], which is oriented towards computer modeling from the moment of its appearance. The purpose of the paper is to describe and test the spectral method for modeling linear filters such as the Butterworth, Linkwitz–Riley, and Chebyshev filters. In this paper, the Linkwitz–Riley filters are considered to show how the spectral form is applied to describe a cascade of linear filters.
The spectral method is an approximate analytical method. It allows the output signal to be represented analytically as a continuous-time function. While this representation is theoretically exact, an approximate representation is used for computer modeling linear filters. The spectral method allows for the choice of a signal representation basis to ensure the desired properties. For example, orthogonal polynomials and trigonometric functions guarantee smoothness, or differentiability, as well as high approximation accuracy. Piecewise constant functions as a basis should produce results comparable to methods based on time discretization.
The proposed technique reduces filtering to a linear transform of expansion coefficients of both deterministic and random signals relative to the chosen basis, utilizing only algebraic operations. Thus, the spectral method combines the advantages of analytical methods and the simple implementation inherent in numerical methods. To the best of our knowledge, the spectral form of mathematical description has not been previously applied for computer modeling linear filters, in particular the Butterworth, Linkwitz–Riley, and Chebyshev filters.
Traditional methods based on time discretization have some drawbacks [28]. For example, the impulse invariance method suffers from aliasing, which occurs if the discretization frequency is insufficient. The reason is that the information about signal is used only at discrete time moments, i.e., on a null set. In contrast, the spectral method uses the information about signal at a given time interval, and the information for discrete time moments of time is not significant. Another discretization method based on a bilinear transform suffers from frequency warping, which arises from the nonlinear nature of the frequency axis transform. The spectral method does not use any nonlinear transform of the frequency axis and, therefore, does not imply frequency warping.
@@ 2. Description of Linear Filters and Problem Statement
In this section and below, we use standard notations and terms from real and complex analysis [34, 35]. In particular, $\mathds{R}$ denotes the set of real numbers, $\mathds{C}$ represents the set of complex numbers (the complex plane); $\mathrm{i}$ is the imaginary unit; $\mathop{\mathrm{Re}}z$ and $\mathop{\mathrm{Im}}z$ denote the real and imaginary parts of a number $z\in\mathds{C}$ which give the Cartesian coordinates of a point in the complex plane; $|z|$ and $\arg z$ are its absolute value and argument which determine the polar coordinates of a point on the complex plane; $\bar{z}$ is the complex conjugate of $z$. The sets $\{z\colon\mathop{\mathrm{Re}}z<0\}$ and $\{z\colon\mathop{\mathrm{Im}}z>0\}$ are the left and upper half-planes, respectively. In addition, some definitions and notations from the theory of function spaces are used, namely spaces of square-integrable functions $L_{2}(\mathds{T})$} and $L_{2}(\mathds{T}^{2})$}, as well as their bases [36].
@@ 1. Description by linear differential equations:
where $g(t)$ and $x(t)$ are input and output signals, respectively, $a_{0},a_{1},\dots,a_{n}\in\mathds{R}$ and $b_{0},b_{1},\dots,b_{m}\in\mathds{R}$ are given values ($a_{n}\neq 0$ and $b_{m}\neq 0$), $n$ represents the filter order. The initial conditions are assumed to be zero.
@@ 4. Description by spectral transforms:
The matrix $W$ can be expressed through coefficients $a_{0},a_{1},\dots,a_{n}$ and $b_{0},b_{1},\dots,b_{m}$:
@@
A linear filter can be interpreted as a linear operator defined on the signal space. In explicit form, this linear operator is specified by the formula (2), and the impulse response function $k(t,\tau)$ defines its kernel. The two-dimensional non-stationary transfer function $W$ gives the spectral, or matrix, representation of such a linear operator relative to the signal space basis. In this paper, the impulse response function can be considered as a function of one variable, namely $k(\eta)$ for $\eta=t-\tau$, and it corresponds to the transfer function $H(s)$.
@@ 3. The Spectral Form of Mathematical Description
1. The term “non-stationary” is used when describing signals and linear systems in the spectral form. In [31], as well as in earlier and later publications on the spectral theory of control systems, the non-stationarity is understood in the following sense. Boundaries of the time segment, on which a linear system should be analyzed, change over time (both boundaries may change, or only one of them, usually the right boundary), i.e., the time segment is non-stationary. In this paper, the time segment is assumed to be fixed, or stationary.
2. Modern publications on the spectral form of mathematical description use different terminology. For example, in [37, 38], the presentation is in terms of spectral characteristics of functions and linear operators. However, the results presented in this paper pertain to signal processing, and in this context, the classical terminology from [31] is more preferable.
3. Although when describing linear systems by integral and spectral transforms, signals and a linear system itself are represented differently (in the first case, as functions of a complex variable; in the second case, as infinite column matrices and infinite matrices), these forms exhibit similar properties. Both forms facilitate the transition from differential and integral relations to algebraic ones between input and output signals. Furthermore, for both forms, characteristics of a linear system are easily expressed through coefficients of the left-hand and right-hand sides of the equation (1).
4. The formulae (4) and (7) have the same structure that enables to easily express the two-dimensional non-stationary transfer function by the transfer function and vice versa. In addition, the representation of the transfer function
$H(s)=\frac{b_{m}(s-\nu_{1})(s-\nu_{2})\ldots(s-\nu_{m})}{a_{n}(s-s_{1})(s-s_{2})\ldots(s-s_{n})},$ (8)
@@
$W=\frac{b_{m}}{a_{n}}\,(P-s_{1}E)^{-1}(P-s_{2}E)^{-1}\ldots(P-s_{n}E)^{-1}(P-\nu_{1}E)(P-\nu_{2}E)\ldots(P-\nu_{m}E).$ (9)
5. The domain for the spectral transform of functions of one variable is the space $L_{2}(\mathds{T})$}. However, the spectral transform can also be applied to generalized functions [31], as well as random processes, including generalized random processes of the white noise type [33]. For the listed variants, the result of the spectral transform is the non-stationary spectral characteristic. The domain for the spectral transform of functions of two variables is also not limited to the space $L_{2}(\mathds{T}^{2})$}, and generalized functions belong to such a domain. Four of the simplest linear systems are considered below, and only one of them corresponds to the impulse transfer function from $L_{2}(\mathds{T}^{2})$}. Impulse transfer functions can be the generalized functions, and two-dimensional non-stationary transfer functions are defined for them.
6. When analyzing and synthesizing linear filters, frequency characteristics determined by the Fourier transform are used first. The spectral transform is associated with the generalized Fourier series, and arbitrary orthonormal bases can be used to represent signals. Below, trigonometric functions are used as a basis, but other bases can be used with the same success, e.g., orthogonal polynomials, piecewise constant functions, etc.
@@
The elements of the matrix $A$ (for the proportional block with gain $a(t)$) are calculated using the formula
$A_{ij}=\int_{\mathds{T}}a(t)q(i,t)q(j,t)dt,\ \ \ i,j=0,1,2,\dots$
For elements of matrices $P^{-1}$ and $P$ (for integral and derivative blocks), the following relations are satisfied:
@@
where $\tau<0$ corresponds to a time delay, and $\tau>0$ corresponds to a time advance. As a rule, in the latter formula, basis functions $q(i,t)$ are extended by zero: $q(i,t+\tau)=0$ for $t+\tau\notin\mathds{T}$, $i=0,1,2,\dots$ However, we redefine them differently (further, it is described in more detail). In the problem under consideration, this improves the quality of modeling linear filters in examples.
@@
forming the orthonormal basis of $L_{2}(\mathds{T})$}, are offered below. In these formulae, the indices take the following values: $i,j=1,2,\dots$ and $i\neq j$.
@@
Formally, the domain of basis functions is the segment $\mathds{T}$. However, basis functions can be defined beyond $\mathds{T}$ in a natural way, i.e., values of basis functions beyond $\mathds{T}$ are determined by the same formula as on $\mathds{T}$. Such a definition is used for calculating elements of the matrix $S$, and another method is discussed in Appendix.
@@
The elements of the column matrix $F$ corresponding to the signal $f(t)=\sin\omega t$ are
$F_{0}=\frac{1-\cos T\omega}{\sqrt{T}\omega},\ \ \ F_{i}=\frac{T\sqrt{2T}\omega[(-1)^{i}\cos T\omega-1]}{i^{2}\pi^{2}-T^{2}\omega^{2}}\ \ \ \text{for}\ \ \ i\pi\neq T\omega,\ \ \ F_{i}=0\ \ \ \text{for}\ \ \ i\pi=T\omega.$
For elements of the column matrix $F$ corresponding to the signal $f(t)=\cos\omega t$, we have
@@ 4.1. The Butterworth Filters
The domain coloring technique [35] can be used to visualize the characteristic polynomial of the Butterworth filter (the transfer function denominator). Here, a version of this technique based on the HLS color model is applied, i.e.,
@@
This technique assumes that the value of a function of a complex variable is coded by color; namely, hue and lightness correspond to argument and absolute value, while saturation takes the maximum value. It allows one to plot the graph of the function $f(z)\colon\mathds{C}\to\mathds{C}$ on a plane, which corresponds to the function $\tilde{f}\colon\mathds{R}^{2}\to\mathds{R}^{2}$. Converting colors to grayscale preserves information about the absolute value and, consequently, about zeros of the function (zeros of the function corresponds to black dots on the graph). However, information about the argument will be lost.
Figure 1 shows a fragment of the complex plane $\{z\colon\mathop{\mathrm{Re}}z,\mathop{\mathrm{Im}}z\in[-2,2]\}$, the filter order is $n=3$. Zeros of the characteristic polynomial (transfer function poles) correspond to three points on the unit circle: 1-1, $(-1\pm\mathrm{i}\sqrt{3})/2$.
@@ 4.3. The Chebyshev Type I Filters
A visualization for the characteristic polynomial of the Chebyshev Type I filter (the transfer function denominator without the coefficient $\gamma$) is carried out in the same way as for the Butterworth filter; the filter order is $n=5$, see Figure 3. Zeros of the characteristic polynomial (transfer function poles) correspond to five points; they lie on the ellipse that can be obtained by scaling the unit circle along the axes.
@@ 4.4. The Chebyshev Type II Filters
The transfer function of the $n$th-order Chebyshev Type II filter is specified using previously defined parameters (in this case, the parameter $\varepsilon>0$ is the stopband ripple factor), namely
$H^{\scriptscriptstyle\mathrm{CII}}(s)=\frac{\varkappa(s-\nu_{1}^{\scriptscriptstyle\mathrm{CII}})\dots(s-\nu_{n}^{\scriptscriptstyle\mathrm{CII}})}{(s-s_{1}^{\scriptscriptstyle\mathrm{CII}})\dots(s-s_{n}^{\scriptscriptstyle\mathrm{CII}})}=\varkappa\biggl[\,\prod\limits_{k=1}^{n}(s-s_{k}^{\scriptscriptstyle\mathrm{CII}})\biggr]^{-1}\prod\limits_{\begin{smallmatrix}k=1\\ 2k-1\neq n\end{smallmatrix}}^{n}(s-\nu_{k}^{\scriptscriptstyle\mathrm{CII}}),$ (23)
@@
$\nu_{k}^{\scriptscriptstyle\mathrm{CII}}=\frac{\mathrm{i}}{B_{k}}=\frac{\mathrm{i}}{\mathop{\mathrm{Im}}s_{k}^{\scriptscriptstyle\mathrm{BW}}},\ \ \ s_{k}^{\scriptscriptstyle\mathrm{CII}}=\frac{1}{s_{k}^{\scriptscriptstyle\mathrm{CI}}}=\frac{\bar{s}_{k}^{\scriptscriptstyle\mathrm{CI}}}{|s_{k}^{\scriptscriptstyle\mathrm{CI}}|^{2}},\ \ \ k=1,\dots,n,$ (24)
and $\varkappa=n^{n\,(\mathrm{mod}\,2)}\varepsilon$ is the normalizing coefficient ($\varkappa=\varepsilon$ for even $n$ and $\varkappa=n\varepsilon$ for odd $n$).
The product $(s-\nu_{1}^{\scriptscriptstyle\mathrm{CII}})\dots(s-\nu_{n}^{\scriptscriptstyle\mathrm{CII}})$ in the formula (23) does not contain a factor corresponding to $B_{k}=0$, and
$\prod\limits_{\begin{smallmatrix}k=1\\ 2k-1\neq n\end{smallmatrix}}^{n}(s-s_{k}^{\scriptscriptstyle\mathrm{CII}})=1$
for $n=1$ (a similar remark holds to other formulae in this subsection).
The transfer function zeros are roots of the equation $T_{n}(\mathrm{i}/\nu)=0$, which form complex conjugate pairs. They are expressed through zeros of the Chebyshev polynomial by the inversion and the coefficient $\mathrm{i}$. The transfer function poles can be defined as roots of the equation $\varepsilon^{2}T_{n}^{2}(1/z)=-1$, belonging to the upper half-plane, with the coefficient $\mathrm{i}$. Using relations (11), (19), and (24), we have
@@
$W^{\scriptscriptstyle\mathrm{CII}}(P)$ $=\varkappa\bigl((P-s_{1}^{\scriptscriptstyle\mathrm{CII}}E)\dots(P-s_{n}^{\scriptscriptstyle\mathrm{CII}}E)\bigr)^{-1}(P-\nu_{1}^{\scriptscriptstyle\mathrm{CII}}E)\dots(P-\nu_{n}^{\scriptscriptstyle\mathrm{CII}}E)$ (26)
$=\varkappa\biggl[\,\prod\limits_{k=1}^{n}(P-s_{k}^{\scriptscriptstyle\mathrm{CII}}E)\biggr]^{-1}\prod\limits_{\begin{smallmatrix}k=1\\ 2k-1\neq n\end{smallmatrix}}^{n}(P-\nu_{k}^{\scriptscriptstyle\mathrm{CII}}E).$
@@
and $(s-s_{k}^{\scriptscriptstyle\mathrm{CII}})(s-s_{n-k+1}^{\scriptscriptstyle\mathrm{CII}})=s^{2}+2\alpha A_{k}s/\rho_{k}^{2}+1/\rho_{k}^{2}$. Additionally, $(s-\nu_{k}^{\scriptscriptstyle\mathrm{CII}})(s-\nu_{n-k+1}^{\scriptscriptstyle\mathrm{CII}})=s^{2}+1/|B_{k}|^{2}$. This yields the expression equivalent to the formula (23):
$H^{\scriptscriptstyle\mathrm{CII}}(s)=\left\{\begin{array}[]{ll}\displaystyle\varkappa\biggl[\,\prod\limits_{k=1}^{n/2}\biggl(s^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,s+\frac{1}{\rho_{k}^{2}}\biggr)\biggr]^{-1}~\prod\limits_{k=1}^{n/2}\biggl(s^{2}+\frac{1}{|B_{k}|^{2}}\biggr)&\text{for even $n$},\\ \displaystyle\varkappa\biggl[\biggl(s+\frac{1}{\alpha}\biggr)\prod\limits_{k=1}^{(n-1)/2}\biggl(s^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,s+\frac{1}{\rho_{k}^{2}}\biggr)\biggr]^{-1}~\prod\limits_{k=1}^{(n-1)/2}\biggl(s^{2}+\frac{1}{|B_{k}|^{2}}\biggr)&\text{for odd $n$},\end{array}\right$. (27)
@@
$W^{\scriptscriptstyle\mathrm{CII}}(P)$ (28)
$=\left\{\begin{array}[]{ll}\displaystyle\varkappa\biggl[\,\prod\limits_{k=1}^{n/2}\biggl(P^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,P+\frac{1}{\rho_{k}^{2}}\,E\biggr)\biggr]^{-1}~\prod\limits_{k=1}^{n/2}\biggl(P^{2}+\frac{1}{|B_{k}|^{2}}\,E\biggr)&\text{for even $n$},\\ \displaystyle\varkappa\biggl[\biggl(P+\frac{1}{\alpha}\,E\biggr)\prod\limits_{k=1}^{(n-1)/2}\biggl(P^{2}+\frac{2\alpha A_{k}}{\rho_{k}^{2}}\,P+\frac{1}{\rho_{k}^{2}}\,E\biggr)\biggr]^{-1}~\prod\limits_{k=1}^{(n-1)/2}\biggl(P^{2}+\frac{1}{|B_{k}|^{2}}\,E\biggr)&\text{for odd $n$}.\end{array}\right$.